# Many-Body Localization in the Age of Classical Computing

### v2

@@ V.3 Tensor network approaches
The tensor network representations of many-body states enable efficient time evolution algorithms [710] such as Time Evolving Block Decimation (TEBD) [676, 677], which is essentially equivalent to time-dependent density matrix renormalization group algorithm [706, 149], or its consecutive improvement, the Time-Dependent Variational Principle (TDVP) [259, 332, 260, 457, 246]. Starting from a given state represented in a matrix product form those algorithms produce an approximation of the time evolved state |$\ket{\psi_{\chi}(t)}$ for a given bond dimension $\chi$. Careful monitoring of the changes of the state |$\ket{\psi_{\chi}(t)}$ with the bond dimension $\chi$ allows to verify whether the results of such simulations are converged with $\chi$. The convergence occurs when $\chi$ is sufficiently large and further increase of $\chi$ does not lead to appreciable changes of the quantities of interest. The tensor network algorithms allow to compute time evolution of many-body systems comprising of hundreds of lattice sites in one-dimension [601, 174, 331, 726, 110, 111] and even to tackle time evolution of quasi-1D or small two-dimensional systems [171].
@@
The growth of entanglement under coherent evolution of interacting many-body systems is the main limiting factor of those methods. The bigger the entanglement entropy, the larger is the value of $\chi$ needed to faithfully represent the time evolved state. For instance, if the entanglement entropy follows a volume-law and scales proportionally to the subsystem size of a one-dimensional system, the required $\chi$ scales exponentially with $L$, implying that there is essentially no gain over the exact representation of the state as a vector in the $\mathcal{N}$ dimensional Hilbert space. This, in turn, means that for a given bond dimension $\chi$ there exists a time $t_{\chi}$ such that for $t<t_{\chi}$ the state |$\ket{\psi_{\chi}(t)}$ is, for a given observable and with a desired accuracy, a sufficiently good approximation of |$\ket{\psi(t)}$. The entanglement spreads ballistically [324] in typical chaotic many-body systems, which implies that the time $t_{\chi}$, for bond dimensions realistic for present-day calculations, is of the order of $10$ tunneling times. However, an increase of disorder strength implies emergence of MBL phenomenology and a significant slow-down of the dynamics, which enables to calculate time evolution of systems of few hundreds of lattice sites up to time scales of the order of few thousand tunneling times [604]. This nicely parallels the claim of the slow, logarithmic in time entanglement entropy growth for the MBL-type of dynamics [697, 47].
@@
Light cone renormalization group [195] is a notable approach that allows to study 1D many-body systems directly in the thermodynamic limit $L\to\infty$. This method relies on the fact that the unitary time evolution of an operator can be restricted to an effective light cone due to the Lieb-Robinson bounds. The required translational invariance can be assured by enlarging the on-site Hilbert space and considering discrete disorder models, see also Sec. X.3, enabling studies of MBL directly in the thermodynamic limit [28, 194]. Discrete disorder models allow for using of numerical linked cluster expansion [521] for studies of quench experiments in ergodic and MBL regimes in the $L\to\infty$ limit [639].
@@
Another type of tensor-network based approaches to MBL was opened by [479, 490], who proposed that the full Hamiltonian of an MBL system can be represented as a low-depth quantum circuit. This approach was demonstrated to be useful for investigation of two-dimensional systems at experimentally relevant time scales [699, 700, 675] and even enables to make predictions about three-dimensional systems [124]. However, the range of time scales and system sizes at which those methods are reliable is yet to be understood.
@@ V.4 Outlook
A typical quench scenario in which a many-body system is initialized in a certain out-of-equilibrium state and evolves under Hamiltonian $H$, can be investigated with full ED. However, the sparsity of many-body Hamiltonians makes Chebyshev and Krylov time propagation schemes much more efficient tools for computation of time evolution of many-body systems.
The ED methods and Chebyshev/Krylov time propagation schemes provide numerically exact results, which makes them especially useful in studies of the crossover between ETH and MBL regimes. Hence, in the following sections, we will be focusing on the results obtained with those methods. Nevertheless, the tensor network approaches enable access to much larger system sizes and, provided that their convergence is carefully monitored, yield very useful insights into thermalization of many-body systems.
Finally, we would like to briefly touch upon other numerical methods relevant for studies of ETH and MBL. Algorithms aimed at a direct calculation of LIOMs in MBL systems include approaches based on time evolution of local operators [117, 418], displacement operators [516, 455], flow equations [480, 651, 308, 653, 654, 383], or direct numerical optimization [450, 481, 360, 6]. Ref. [650] argues that the flow equation techniques combined with a method of scrambling transforms yield a method for exploration of intermediate-scale time evolution in many-body systems. Microcanonical Lanczos method (MCLM) introduced in [295] combines Lanczos algorithm and random sampling to investigate finite-temperature static and dynamical quantities of many-body systems and can used to study transport properties, see e.g. [419, 49]. The transport properties can be also studied in an open system approach, involving the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation [243, 375] which can be simulated with tensor network algorithm that calculates the non-equilibrium steady state of open the system [510].
Recently, machine learning approaches have found new applications to quantum sciences [155], see also earlier [411, 106]. While algorithms for an automated distinction between the ETH and MBL behaviors were proposed [287, 174], a thorough analysis [647] indicates a large sensitivity of the quantitative results for the ETH-MBL transition with respect to the network structure and training parameters, preventing one from reaching a definite conclusions about the transition. Nevertheless, machine learning approaches can be useful for tasks such as the analysis of details of time evolution of disordered systems [632, 339] or classification of families of non-ergodic eigenstates of many-body Hamiltonians [633]. In addition, unsupervised machine learning approaches have been successful in identifying phases of quantum many-body systems [527, 413, 281] and have been recently employed in the context of Anderson transition [672].
@@ VI.1 Spectral statistics
As postulated by Wigner, Dyson, Porter and other luminaries of random matrix theory, the statistical properties of the energy levels of complex systems are well approximated by random matrices [410, 258]. Quantum chaotic systems exhibit level repulsion [80], with the probability to have a small level spacing (compared to the average level spacing) being suppressed differently depending only on the global symmetry of the Hamiltonian. For example, the Gaussian Orthogonal Ensemble (GOE) applies to time-reversal invariant systems such as (7), see [410, 258]. In contrast, systems in the MBL regime, whose Hamiltonians in finite systems commute with the set of LIOMs [^12] The LIOMs, i.e., the quasi-local operators which commute with the Hamiltonian, can be numerically constructed in the MBL regime at finite $L$, irrespective of their status in the $L\to\infty$ limit. (9), have uncorrelated energy levels with Poisson statistics. The ease of accessing the eigenvalues $E_{n}$ through exact diagonalization studies of $\hat{H}$, coupled with the clear distinction between ergodic and MBL systems, makes investigating spectral statistics a natural choice in numerical studies of MBL.
@@ VI.1.1 Gap ratio
While earlier studies considered level spacing statistics, they suffered from the procedure of level unfolding (taking into account vastly different density of states in various physical models) [258]. That procedure is avoided by considering a gap ratio, introduced in [451]. The gap ratio is a dimensionless quantity defined as
@@
where $\delta_{n}=E_{n+1}-E_{n}$ is the gap between two adjacent eigenvalues in the spectrum of the system. Studies of MBL typically concern the middle of the spectrum, where the density of states is the largest. Averaging $r_{n}$ over eigenvalues in this energy range and over different disorder realizations yields an average gap ratio, denoted by $\overline{r}$. Comparing the average gap ratio, $\overline{r}$, to the values $\overline{r}_{GOE}\approx 0.5307$ and $\overline{r}_{PS}\approx 0.3863$, which are characteristic for the GOE [36] and Poisson level statistics, respectively, is for most physical systems a way of assessing whether the system is ergodic or not [^13] This remark concerns a generic case. However, the lack of ergodicity associated with the existence of a particular class of eigenstates, for instance with quantum scar states [669], may not be detectable with the probes of the level statistics..
The average gap ratio, $\overline{r}$, calculated for disordered many-body systems is the prototypical quantity featuring the trends illustrated in Fig. 2(b). Here, we present results for the $J_{1}$-$J_{2}$ model with the Hamiltonian
@@
where $\hat{H}_{\textit{XXZ}}$ is the Hamiltonian of the disordered XXZ spin chain, see Eq. (7), $\Delta=0.55$, and periodic boundary conditions are assumed. In contrast to the disordered XXZ spin chain, however, the limit $W\to 0$ is not Bethe ansatz integrable, which simplifies the analysis in the weak disorder limit. The Hamiltonian $\hat{H}_{J_{1}-J_{2}}$, similarly to $\hat{H}_{\textit{XXZ}}$, conserves the $z$ component of total spin, $\hat{S}^{z}=\sum_{i=1}^{L}\hat{S}^{z}_{i}$. We follow the standard choice in the literature on MBL and consider the sector with $\hat{S}^{z}=0$, which is the largest symmetry sector of the model141414 A meaningful analysis of level statistics is possible only when the global symmetries of the Hamiltonian are resolved. Before the resolution of the symmetries, the spectrum of an ergodic system is a superposition of independent GOE/GUE spectra with statistical properties approaching the Poisson case with the increase of the number of symmetry sectors, see [232] for a recent discussion of this well known problem. . Moreover, the results depend on the energy density at which the gap ratio is calculated [393]. In the following, to extract $\overline{r}$, we average (26) over less than $5\%$ of energies in the middle of the spectrum of the system. The average gap ratio $\overline{r}$ for the $J_{1}$-$J_{2}$ model is shown in Fig. 4(a). While we here present the concrete example of the $J_{1}$-$J_{2}$ model, analogous trends are found in the results for XXZ spin chain and other disordered many-body systems, see Sec. X.1. Note, however, that the quantitative values of disorder strengths in the crossover regime are larger in the $J_{1}$-$J_{2}$ model compared to the XXZ model.
@@
Figure 4: ETH-MBL crossover in the $J_{1}$-$J_{2}$ spin chain. The average gap ratio $\overline{r}$ for system size $L$, shown in (a), changes upon a variation of disorder strength $W$ between the GOE value $\overline{r}_{GOE}\approx 0.5307$ in the ETH regime and Poisson value $\overline{r}_{PS}\approx 0.3863$ in the MBL regime. The crossing point of $\overline{r}(W)$ curves, denoted in red, drifts with increasing disorder strength. Panel (b) illustrates the extraction of the $L$ dependent crossing point $W^{*}(L)$ for $L=14$ and $\Delta L=2$. Panel (c) shows how to extract the disorder strength $W^{T}(L)$ at which system creases to follow the ETH prediction.
@@
For a sufficiently weak disorder in the $J_{1}$-$J_{2}$ model, $W\lessapprox 6$, there is a clear trend towards $\overline{r}_{GOE}$ predicted for a system that follows ETH. In contrast, for the strong disorder, $W\gtrapprox 10$, we observe a monotonous with $L$ decrease of $\overline{r}$ towards the Poisson value $\overline{r}_{PS}$, consistent with an emergent localization of the system. While the slope of the curves $\overline{r}(W)$ in the intermediate region increases with system size, the results do not follow the simple-to-interpret trend illustrated in Fig. 2(a). On the contrary, there is a significant drift of the crossing point of the $\overline{r}(W)$ curves. While this problem was noted in the early work [451], the authors observed that the drift of the crossing point slows down with the increase of the system size, providing a hint of the presence of the MBL phase. The presence of the MBL phase, based on a more detailed study of the XXZ spin chain [460], became for over a decade the prevailing interpretation of the results for strongly disordered many-body systems. One of the deficiencies of this interpretation was highlighted in [690], where it was shown that the results for different $L$ collapse upon a rescaling $\overline{r}(W)\to\overline{r}(W/L)$ close to the ETH regime, i.e., for $W$ at which $\overline{r}(W)>\overline{r}_{GOE}-\Delta\overline{r}$ (where $\Delta\overline{r}\ll 1$ is a constant). The latter rescaling, taken literally, implies that the value of $\overline{r}$ remains fixed if the disorder strength increases proportionally to the system size $L$, suggesting that there is no MBL phase and $\overline{r}(W)\to\overline{r}_{GOE}$ at any $W$ for $L\to\infty$.
@@
Finally, it is possible that neither of the scalings (28), (29), even after modifications including the putative sub-leading terms, persists in the $L\to\infty$ limit. This can never be entirely excluded solely based on analysis of numerical data at finite $L$. Nevertheless, the numerical results presented in Fig. 4(a) do not provide any substantial premises suggesting quantitative changes of the trends at larger system sizes that are out of the reach of present-day classical supercomputers. This leaves out scenarios 1. and 2., which cannot be unambiguously differentiated by the analysis of presently available numerical results. Scenario 1. implies the existence of an MBL phase and a transition between ergodic and MBL phases at a finite disorder strength. (For the disordered XXZ spin chain, an extrapolation of the scaling (29) to $L\to\infty$ limit yields $W_{C}=W_{\infty}\approx 5.4$, larger than $W_{C}\approx 3.7$ proposed in [393].) In contrast, scenario 2. predicts that there is no MBL phase and that the disordered spin chains become ergodic at any $W$ if the system size is sufficiently large.
While the above perspective on the numerical results for the disordered spin chains appears grim, it may be possible to distinguish scenarios 1. and 2. in the not-so-distant future. Extrapolation of the scalings (29) and (28) shows that they become incompatible at system size $L_{0}\approx 50$ for the disordered XXZ spin chain. The exact calculation of eigenstates in the middle of the spectrum for $L=50$ is far beyond the capabilities of present-day supercomputers. However, the signatures of the breakdown of (29) or of (28) may be observable at significantly smaller system sizes, only slightly beyond $L$ studied so far.
In passing, let us note that a crossover between delocalized and localized phases of the Anderson model on random regular graphs shows phenomenology similar to the ETH-MBL crossover [660, 661]. Analysis of the crossover with disorder strengths $W^{T}(L)$ and $W^{*}(L)$ shows a breakdown of the linear scaling of $W^{T}(L)$ and indicates that extrapolation of the $W^{*}(L)$ allows to reproduce accurately the critical disorder strength for Anderson localization transition on random regular graphs [595]. While those results provide a non-trivial test of the method of analysis of the delocalized-localized crossover with disorder strengths $W^{T}(L)$ and $W^{*}(L)$, the Anderson model on random regular graphs, as a single-particle model, is crucially different from the problem of MBL in disordered many-body systems. Finally, we would like to note that the deviations from the linear shift of $W^{T}(L)$ with $L$, which constitute another premise of the existence of an MBL phase in the thermodynamic limit, were found in disordered Floquet models [596], see Sec. X.4.

### v3

@@ V.3 Tensor network and other approaches
The tensor network representations of many-body states enable efficient time evolution algorithms [766] such as Time Evolving Block Decimation (TEBD) [731, 732], which is essentially equivalent to time-dependent density matrix renormalization group algorithm [762, 155], or its consecutive improvement, the Time-Dependent Variational Principle (TDVP) [275, 360, 276, 495, 262]. Starting from a given state represented in a matrix product form, those algorithms produce an approximation of the time evolved state |$\ket{\psi_{\chi}(t)}$ for a given bond dimension $\chi$. Careful monitoring of the changes of the state |$\ket{\psi_{\chi}(t)}$ with the bond dimension $\chi$ allows to verify whether the results of such simulations are converged with $\chi$. The convergence occurs when $\chi$ is sufficiently large and a further increase of $\chi$ does not lead to appreciable changes of the quantities of interest. The tensor network algorithms allow to compute time evolution of many-body systems comprising of hundreds of lattice sites in one dimension [650, 183, 359, 782, 114, 115] and even to tackle time evolution of quasi-1D or small two-dimensional systems [180].
@@
The growth of entanglement under coherent evolution of interacting many-body systems is the main limiting factor of those methods. The bigger the entanglement entropy, the larger are the values of $\chi$ needed to faithfully represent the time evolved state. For instance, if the entanglement entropy follows a volume-law and scales proportionally to the subsystem size of a one-dimensional system, the required $\chi$ scales exponentially with $L$, implying that there is essentially no gain over the exact representation of the state as a vector in the $\mathcal{N}$ dimensional Hilbert space. This, in turn, means that for a given bond dimension $\chi$ there exists a time $t_{\chi}$ such that for $t<t_{\chi}$, the state |$\ket{\psi_{\chi}(t)}$ is, for a given observable and with a desired accuracy, a sufficiently good approximation of |$\ket{\psi(t)}$. The entanglement spreads ballistically [350] in typical chaotic many-body systems, which implies that the time $t_{\chi}$, for bond dimensions realistic for present-day calculations, is of the order of $10$ tunneling times. However, an increase of disorder strength implies emergence of MBL phenomenology and a significant slow-down of the dynamics, which enables to calculate time evolution of systems of few hundreds of lattice sites up to time scales of the order of few thousand tunneling times [653]. This nicely parallels the claim of the slow, logarithmic in time entanglement entropy growth for the MBL-type of dynamics [753, 50].
@@
Light cone renormalization group [205] is a notable approach that allows to study 1D many-body systems directly in the thermodynamic limit $L\to\infty$. This method relies on the fact that the unitary time evolution of an operator can be restricted to an effective light cone due to the Lieb-Robinson bounds [407]. The required translational invariance can be assured by enlarging the on-site Hilbert space and considering discrete disorder models, see also Sec. X.3, enabling studies of MBL directly in the thermodynamic limit [30, 204]. Discrete disorder models allow for using of numerical linked cluster expansion [565] for studies of quench experiments in ergodic and MBL regimes in the $L\to\infty$ limit [690].
@@
Another type of tensor-network based approaches to MBL was opened by [517, 529], who proposed that the full Hamiltonian of an MBL system can be represented as a low-depth quantum circuit. This approach was demonstrated to be useful for investigation of two-dimensional systems at experimentally relevant time scales [755, 756, 730, 403] and even enables to make predictions about three-dimensional systems [130]. However, the range of time scales and system sizes at which those methods are reliable is yet to be understood.
@@ V.4 Outlook
A typical quench scenario in which a many-body system is initialized in a certain out-of-equilibrium state and evolves under Hamiltonian $\hat{H}$, can be investigated with full ED. However, the sparsity of many-body Hamiltonians makes Chebyshev and Krylov time propagation schemes much more efficient tools for computation of time evolution of many-body systems.
The ED methods and Chebyshev/Krylov time propagation schemes provide numerically exact results, which makes them especially useful in studies of the crossover between ETH and MBL regimes. Hence, in the following sections, we will focus on the results obtained with those methods. Nevertheless, the tensor network approaches enable access to much larger system sizes and, provided that their convergence is carefully monitored, yield very useful insights into thermalization of many-body systems.
Finally, we would like to briefly touch upon other numerical methods relevant for studies of ETH and MBL. Algorithms aimed at a direct calculation of LIOMs in MBL systems include approaches based on time evolution of local operators [122, 454], displacement operators [557, 493], flow equations [518, 703, 334, 705, 706, 419], or direct numerical optimization [488, 519, 390, 6]. It was argued in [702] that the flow equation techniques combined with a method of scrambling transforms yield a method for exploration of intermediate-scale time evolution in many-body systems. Microcanonical Lanczos method (MCLM) introduced in [320] combines Lanczos algorithm and random sampling to investigate finite-temperature static and dynamical quantities of many-body systems and can be used to study transport properties, see, e.g., [455, 52]. The transport properties can also be studied in an open system approach, involving the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation [259, 410], which can be simulated with tensor network algorithm that calculates the non-equilibrium steady state of the open system [551].
Recently, machine learning approaches have found new applications to quantum sciences [161], see also earlier works [447, 109]. While algorithms for an automated distinction between the ETH and MBL behaviors were proposed [308, 183], a thorough analysis [699] indicates a large sensitivity of the quantitative results for the ETH-MBL transition with respect to the network structure and training parameters, preventing one from reaching a definite conclusions about the transition. Nevertheless, machine learning approaches can be useful for tasks such as the analysis of details of time evolution of disordered systems [683, 368] or classification of families of non-ergodic eigenstates of many-body Hamiltonians [684]. In addition, unsupervised machine learning approaches have been successful in identifying phases of quantum many-body systems [572, 449, 300] and have been recently employed in the context of Anderson transition [726].
@@ VI.1 Spectral statistics
As postulated by Wigner, Dyson, Porter and other pioneers of random matrix theory, the statistical properties of the energy levels of complex systems are well approximated by random matrices [446, 274]. Quantum chaotic systems exhibit level repulsion [83], with the probability distribution of level spacing depending only on the global symmetries of the Hamiltonian. For example, the Gaussian Orthogonal Ensemble (GOE) applies to time-reversal invariant systems such as (7), see [446, 274]. In contrast, systems in the MBL regime, whose Hamiltonians in finite systems commute with the set of LIOMs [^20] The LIOMs, i.e., the quasi-local operators which commute with the Hamiltonian, can be numerically constructed in the MBL regime at finite $L$, irrespective of their status in the $L\to\infty$ limit. (9), have uncorrelated energy levels with Poisson statistics. The ease of accessing the eigenvalues $E_{n}$ through exact diagonalization studies of $\hat{H}$, coupled with the clear distinction between ergodic and MBL systems, makes investigating spectral statistics a natural choice in numerical studies of MBL.
@@ VI.1.1 Gap ratio
While earlier studies considered level spacing statistics, they suffered from the procedure of level unfolding [274]. That procedure is avoided considering a gap ratio, introduced in [489]. The gap ratio is a dimensionless quantity defined as
@@
where $\delta_{n}=E_{n+1}-E_{n}$ is the gap between two adjacent eigenvalues in the spectrum of the system. Studies of MBL typically concern the middle of the spectrum, where the density of states is the largest. Averaging $r_{n}$ over eigenvalues in this energy range and over different disorder realizations yields an average gap ratio, denoted by $\overline{r}$. Comparing the average gap ratio, $\overline{r}$, to the values $\overline{r}_{GOE}\approx 0.5307$ and $\overline{r}_{PS}=2\ln 2-1\approx 0.3863$, which are characteristic for the GOE [38] and Poisson level statistics, respectively, is for most physical systems a way of assessing whether the system is ergodic or not [^21] This remark concerns a generic case. However, the lack of ergodicity associated with the existence of a particular class of eigenstates, for instance with quantum scar states [723], may not be detectable with the probes of the level statistics..
The average gap ratio, $\overline{r}$, calculated for disordered many-body systems, is the prototypical quantity featuring the trends illustrated in Fig. 2(b). Here, we present results for the $J_{1}$-$J_{2}$ model with the Hamiltonian
@@
where $\hat{H}_{\textit{XXZ}}$ is the Hamiltonian of the disordered XXZ spin chain, see Eq. (7), $\Delta=0.55$, and periodic boundary conditions are assumed. In contrast to the disordered XXZ spin chain, however, the limit $W\to 0$ is not Bethe ansatz integrable, which simplifies the analysis in the weak disorder limit. The Hamiltonian $\hat{H}_{J_{1}-J_{2}}$, similarly to $\hat{H}_{\textit{XXZ}}$, conserves the $z$ component of total spin, $\hat{S}^{z}=\sum_{i=1}^{L}\hat{S}^{z}_{i}$. We follow the standard choice in the literature on MBL and consider the sector with $\hat{S}^{z}=0$, which is the largest symmetry sector of the model222222 A meaningful analysis of level statistics is possible only when the global symmetries of the Hamiltonian are resolved. Before the resolution of the symmetries, the spectrum of an ergodic system is a superposition of independent GOE/GUE spectra with statistical properties approaching the Poisson case with the increase of the number of symmetry sectors, see [248] for a recent discussion of this well known problem. . Moreover, the results depend on the energy density at which the gap ratio is calculated [430]. In the following, to extract $\overline{r}$, we average (26) over less than $5\%$ of energies in the middle of the spectrum of the system. The average gap ratio $\overline{r}$ for the $J_{1}$-$J_{2}$ model is shown in Fig. 4(a). While we here present the concrete example of the $J_{1}$-$J_{2}$ model, analogous trends are found in the results for XXZ spin chain and other disordered many-body systems, see Sec. X.1. Note, however, that the quantitative values of disorder strengths in the crossover regime are larger in the $J_{1}$-$J_{2}$ model compared to the XXZ model.
@@
Figure 4: ETH-MBL crossover in the $J_{1}$-$J_{2}$ spin-1/2 chain. The average gap ratio $\overline{r}$ for system size $L$, shown in (a), changes upon a variation of disorder strength $W$ between the GOE value $\overline{r}_{GOE}\approx 0.5307$ in the ETH regime and Poisson value $\overline{r}_{PS}\approx 0.3863$ in the MBL regime. The crossing point of $\overline{r}(W)$ curves, denoted in red, drifts with increasing disorder strength. Panel (b) illustrates the extraction of the $L$ dependent crossing point $W^{*}(L)$ for $L=14$ and $\Delta L=2$. Panel (c) shows how to extract the disorder strength $W^{T}(L)$ at which system creases to follow the ETH prediction.
@@
For a sufficiently weak disorder in the $J_{1}$-$J_{2}$ model, $W\lessapprox 6$, there is a clear trend towards $\overline{r}_{GOE}$ predicted for a system that follows ETH. In contrast, for the strong disorder, $W\gtrapprox 10$, we observe a monotonous with $L$ decrease of $\overline{r}$ towards the Poisson value $\overline{r}_{PS}$, consistent with an emergent localization of the system. While the slope of the curves $\overline{r}(W)$ in the intermediate region increases with system size, the results do not follow the simple-to-interpret trend illustrated in Fig. 2(a). On the contrary, there is a significant drift of the crossing point of the $\overline{r}(W)$ curves. While this problem was noted in the early paper [489], the authors observed that the drift of the crossing point slows down with the increase of the system size, providing a hint of the presence of the MBL phase. The presence of the MBL phase, based on a more detailed study of the XXZ spin chain [498], became for over a decade the prevailing interpretation of the results for strongly disordered many-body systems. One of the deficiencies of this interpretation was highlighted in [746], where it was shown that the results for different $L$ collapse upon a rescaling $\overline{r}(W)\to\overline{r}(W/L)$ close to the ETH regime, i.e., for $W$ at which $\overline{r}(W)>\overline{r}_{GOE}-\Delta\overline{r}$ (where $\Delta\overline{r}\ll 1$ is a constant). The latter rescaling, taken literally, implies that the value of $\overline{r}$ remains fixed if the disorder strength increases proportionally to the system size $L$, suggesting that there is no MBL phase and $\overline{r}(W)\to\overline{r}_{GOE}$ at any $W$ for $L\to\infty$.
@@
Finally, it is possible that neither of the scalings (28), (29), even after modifications including the putative sub-leading terms, persists in the $L\to\infty$ limit. This can never be entirely excluded solely based on analysis of numerical data at finite $L$. Nevertheless, the numerical results presented in Fig. 4(a) do not provide any substantial premises suggesting quantitative changes of the trends at larger system sizes that are out of the reach of present-day classical supercomputers. This leaves out scenarios 1. and 2., which cannot be unambiguously differentiated by the analysis of presently available numerical results. Scenario 1. implies the existence of an MBL phase and a transition between ergodic and MBL phases at a finite disorder strength. (For the disordered XXZ spin-1/2 chain, an extrapolation of the scaling (29) to $L\to\infty$ limit yields $W_{C}=W_{\infty}\approx 5.4$, larger than $W_{C}\approx 3.7$ proposed in [430].) In contrast, scenario 2. predicts that there is no MBL phase and that the disordered spin chains may become ergodic at any $W$ if the system size is sufficiently large.
While the above perspective on the numerical results for the disordered spin chains appears grim, it may be possible to distinguish Scenarios 1. and 2. in the not-so-distant future. Extrapolation of the scalings (29) and (28) shows that they become incompatible at system size $L_{0}\approx 50$ for the disordered XXZ spin-1/2 chain. The exact calculation of eigenstates in the middle of the spectrum for $L=50$ is far beyond the capabilities of present-day supercomputers. However, the signatures of the breakdown of (29) or of (28) may be observable at significantly smaller system sizes, only slightly beyond $L$ studied so far.
In passing, let us note that a crossover between delocalized and localized phases of the Anderson model on random regular graphs shows phenomenology that shares certain similarities with the ETH-MBL crossover [712, 713]. Analysis of the crossover with disorder strengths $W^{T}(L)$ and $W^{*}(L)$ shows a sub-linear scaling of $W^{T}(L)$ with $L$ and indicates that extrapolation of the $W^{*}(L)$ allows to reproduce accurately the critical disorder strength for Anderson localization transition on random regular graphs [644]. While those results provide a non-trivial test of the method of analysis of the delocalized-localized crossover with disorder strengths $W^{T}(L)$ and $W^{*}(L)$, the Anderson model on random regular graphs, as a single-particle model, is crucially different from the problem of MBL in disordered many-body systems. Finally, we would like to note that the deviations from the linear shift of $W^{T}(L)$ with $L$, which constitute another premise of the existence of an MBL phase in the thermodynamic limit, were found in disordered Floquet models [645], see Sec. X.4.